OpenStudy (anonymous):

Simplify. Express using positive exponents. ***Problem requires exponents. See comments for problem.***

OpenStudy (anonymous):

\[(3y)^{-3}\]

OpenStudy (anonymous):

I'll give you a hint. \[x^{-1 / 2} = {1 \over \sqrt{x}} \]

OpenStudy (anonymous):

That is nothing like what i am doing in math right now

OpenStudy (anonymous):

That is exactly what you are doing. \[x^{-1/2} = {1 \over x^{1/2}}\]

OpenStudy (anonymous):

you flip the fraction. so it the negative exponent is on the top it switches to the bottom and if it is on the bottom it switches to the top. however, you only move the variable/number that has the negative exponent attached to it.